# thanks: These two authors contributed equallythanks: These two authors contributed equally

### v2

@@ Abstract
Quantum emitter-based schemes for the generation of photonic graph states offer a promising, resource efficient methodology for realizing distributed quantum computation and communication protocols on near-term hardware. We present a heralded scheme for making photonic graph states that is compatible with the typically poor photon collection from state-of-the-art coherent quantum emitters. We demonstrate that the construction time for large graph states can be polynomial in the photon collection efficiency, as compared to the exponential scaling of current emitter-based schemes, which assume deterministic photon collection. The additional overhead to achieve this advantage consists of an extra spin system plus one additional spin-spin entangling gate per photon added to the graph. While the proposed scheme enables the generation of graph states for arbitrary applications, we show how it can be further simplified for the specific task of measurement-based computation, leading to significantly higher rates and removing the need for photonic memory in certain computations. As an example use-case of our scheme, we construct a protocol for secure two-party computation that can be implemented efficiently on current hardware. Estimates of the fidelity to produce graph states used in the computation are given, based on current trapped ion experimental benchmarks.
@@ I Introduction
The vast promise of quantum information technologies for faster and more secure computational and information systems relies on entanglement as a primary resource. Traditional gate-based quantum computing requires the ability to perform sequences of joint operations across multiple qubits. An alternative paradigm, known as measurement-based quantum computation (MBQC), realizes universal computation through sequences of single qubit measurements made on an initially prepared entangled resource state [1]. This is an appealing approach for photonic quantum systems as single photon rotations and measurements are straightforward using commercial optical elements, and fast efficient routing solutions are readily available [2]. Furthermore, the sequential nature of photon emission allows entangled states to be built from photons emitted at different times by the same emitter [3]. By using entangled emitters, it then becomes possible to simulate entangling gates between photons and overcome the difficulty of realizing direct photon-photon interactions [4]. Indeed, it is known that computationally useful graph states of photons can be generated using either a small number of coherent quantum emitters [5], a combination of a single emitter and fusion gates [6, 7, 8], or a single quantum emitter in a feedback scheme [9].
@@
All of these aforementioned schemes assume a photon is successfully added to the graph every time an emitter is excited, and we thus term this class of schemes as being “deterministic”. However, efficient collection from individual quantum emitters is a challenge that remains largely out of reach today, making deterministic schemes impractical for generating large graph states on even near-term hardware [10].
We introduce a procedure for building arbitrary photonic graph states using a single emitting spin and a fixed number of auxiliary spins, without any requirement on the collection efficiency of the emitted photons. The rate and size of feasible graph states generated by the proposed method has polynomial dependence on the collection efficiency, compared to the exponential suppression of the feasible graph size for the successive detection of photons required in current deterministic schemes. This enhanced performance is achieved by recognizing the possibility of adding each photon to the graph only after successful spin-photon entanglement generation has been heralded.
In Section II of this work, we describe a non-destructive method for this heralding that is implementable on current hardware by swapping the entanglement to another photon, termed the herald-on-swap (HoS) method. More generally, any non-destructive detection of the photon would be suitable to implement this “emit then add” method of building graphs for arbitrary use. We explain in Section III how we can herald on destructive detection of the photon for certain graph state protocols, including MBQC. This enables much higher production rates and fidelities than any currently feasible non-destructive heralding method. In this case, the emitted photon is measured prior to adding it to the graph, thereby abrogating the need for entanglement swapping or photon storage, and dramatically speeding up graph construction. In this case, the graph state is built only virtually as every photon in the graph is necessarily measured. We refer to this as a herald-on-detection (HoD) method, and Section III provides more detail on when the HoD method can be applied.
In both the HoS and HoD schemes we propose, the photons are emitted over a longer time, when compared to other emitter-based methods for building photonic graph states; this is because most attempts to add a photon to the graph will result in an unsuccessful herald. Consequently, it is the number of excitation attempts that can be made during the coherence time of the emitter that sets the practical limit on the graph size, rather than the collection efficiency. Trapped neutral atoms and ions in vacuum are well suited to this scheme as they exhibit long coherence times ($10^{0}-10^{3}$ seconds) [11, 12] and host high-fidelity deterministic entangling gates ($>99\%$) [13, 14, 15] with moderate collection efficiencies ($\gtrsim 10\%$ with high-NA optics and $\gtrsim 40\%$ with parabolic reflectors) [16, 17, 18, 19]. Higher collection can be achieved by coupling emitters to high quality resonators, but this introduces additional challenges for implementing high-fidelity entangling gates (necessary for generating the higher dimensional graph states required for MBQC) and dramatically reduces the bandwidth of operation [20, 21]. Even higher collection efficiency has been shown with quantum dot sources [22, 7], but in these systems high-fidelity entangling gates between spins are still out of reach and spin coherence times are orders of magnitude shorter than atoms and ions. We also note that some solid-state systems, such as rare-earth emitters, can exhibit excellent coherence, spin-spin entangling gates, and reasonable photon collection via integration into photonic devices [23]. In Section IV we provide estimates of the rate and fidelity for the production of heralded graph states via both of our schemes assuming state-of-the-art trapped ion and photon pair source architectures, which include realistic sources of loss, noise, and decoherence.
Lastly, in Section V we propose a protocol for performing secure two-party classical computation, in which two parties privately compute a function of their input variables with the help of a Referee. Compared to universal quantum computation, our protocol only requires Pauli measurements, and so the entire procedure can be performed using only our HoD method. We demonstrate how secure two-party computation can be achieved using just a single quantum emitter and two auxiliary spin qubits, regardless of the size of the party’s input. We provide fidelity estimates to produce the graph states required for our protocol on current hardware, which suggests this is a highly feasible use-case for quantum networks in the near term. A security analysis is provided to demonstrate how a party can keep their input private. In addition we discuss the resulting bit error rates of the protocol that can be achieved with and without classical error correction.
II Spin-photon entanglement schemes
Figure 1: Herald-on-swap (HoS) scheme. The components are two initial entangled pairs: an emitter spin (red square) entangled with a photon (orange circle) and a signal-idler entangled photon pair (orange and green circle, respectively). We also assume an existing graph $G$ that contains one or more auxiliary spins (black square). Entanglement swapping is attempted between the two initial entangled pairs. Upon success, a two-qubit spin-spin gate deterministically entangles the emitter spin and auxiliary spin. A local complementation operation is then performed on the emitter spin, which is subsequently measured out of the graph. This leaves the idler photon a part of the graph and the emitter spin reinitialized for the next attempted emission.
We first describe the HoS scheme, shown in Figure 1, for making graph states for arbitrary applications. Given the technical challenges associated with photon non-demolition measurement [24], this swapping step is required to herald without destructive measurement of the photonic qubit on near-term hardware. Compared to the deterministic scheme described in [5], we include a local entanglement swapping step between an emitter-photon entangled pair and a pair of entangled photons, and only add the final photon to the graph upon success of that swap. A single emitter is initialized onto an unentangled state and excited to produce a single photon that is entangled with its long-lived internal spin state and collected with some overall efficiency $\eta_{e}$. This can be implemented with a wide variety of quantum emitters including laser-cooled atoms or ions, quantum dots, and defects or dopants in wide-bandgap semiconductors [3, 25, 10, 26, 27]. Photons can be encoded in various degrees of freedom including polarization, time-bin, and frequency [28, 29, 30, 31, 32]. Simultaneously, a pair of entangled photons is probabilistically produced, which can be implemented via standard nonlinear optical processes, such as spontaneous parametric down conversion (SPDC) [33, 34]. One member of the photon pair, the signal photon, is wavelength and bandwidth matched to the emitter photon, and the entangled pair is encoded in the same degree of freedom as the emitter photon. The emitter photon and signal photon are sent to a joint measurement apparatus which, upon a successful measurement outcome, projects the emitter spin and the unmeasured photon, the idler photon, onto an entangled Bell state [35]. Since only certain measurement outcomes correspond to entanglement between the emitter spin and idler photon, and imperfect collection and detection means that fewer than two photons are detected on most attempts, the procedure is repeated until a successful herald is flagged. Following each failed attempt, the emitter is measured and reinitialized for the subsequent attempt at a successful herald.
@@ II Spin-photon entanglement schemes
We assume the emitter is among a set of spin qubits that can be controllably entangled in a pair-wise way via local and deterministic two-qubit spin-spin entangling gates, such as an array of trapped atomic ions or neutral atoms. Upon a successful herald, the emitter spin is deterministically entangled with an auxiliary spin, which may already be part of a larger graph state. The emitter spin is then rotated to a conjugate basis and projectively measured, removing it from the graph without disturbing any of the other qubits. Following a correction to the photon based on the outcome of this measurement, the state of the system is as if the auxiliary spin had directly emitted the photon itself. This “emit then add” method then allows for the construction of arbitrary photonic graph states using the methods described in [5]. Namely, the graph states are built using sequences of two-qubit gates on the spins and local Clifford gates on the photons. The local gates on each photon can all be combined and applied as a single rotation prior to using the graph state in subsequent applications.
The key improvement in our scheme is that the idler photon can be collected with near unit efficiency upon a successful herald [36]. Attaining the improvement of this scheme requires one additional two-qubit spin-spin entangling gate per photon added to the graph plus one additional spin (the emitter), compared to the requirements for generating graphs via known deterministic schemes [37]. In addition, the graph is emitted over a much longer time and thus requires longer spin coherence times. We note an additional benefit of the scheme in that the idler photon can be at virtually any wavelength due to the flexibility of standard entangled photon pair sources [38, 39, 40, 41], which overcomes common challenges due to the inconvenient emission colors of many atomic qubits. We show here that current state-of-the-art trapped ion systems can be designed to produce graph states of tens to hundreds of photons using the proposed method, where extending to thousands would be made possible with reasonable improvements [12, 14, 16].
III A Heralding Scheme for MBQC
@@
The HoS scheme described in the previous section, or any other non-destructive heralding of successful collection of the emitter photon, can be used to build arbitrary photonic graph states. We show in this section that for many applications, including MBQC, non-destructive measurement is not required. For these applications the nodes are measured sequentially with the choice of measurement basis on one node depending on outcomes of previous ones. It is, thus, not necessary to build the full graph state before beginning these measurements [42]. Instead, an emitted photon can be destructively measured as soon as the correct basis for measurement is determined. This measurement thus serves doubly as a prescribed step in the MBQC protocol and as a way for detecting the emitted photon. Upon failure to detect the photon, we reinitialize the emitter and attempt generation again, having left the larger graph undisturbed, as in the HoS scheme. Upon successful detection, we perform all required gates on the emitter and auxiliary spins, measure the emitter out of the graph, and reinitialize the emitter to attempt generation of the next photon in the graph.
@@
This simpler HoD scheme is shown in Figure 2, and it has the advantage that it requires no storage time for the emitted photon prior to measurement. There are limitations for when this scheme can be employed. Namely, a photon can be measured using HoD provided the two conditions are satisfied: (1) the correct measurement basis for that photon, as set by the MBQC protocol, is determined prior to its emission, and (2) this measurement is either Pauli $Z$ or of the form $\cos\phi X+\sin\phi Y$. Condition (1) can be met in general, as the emission order can be chosen to match the measurement order of the computation, albeit at the cost of using extra auxiliary spins and two-qubit gates in some cases [37]. The second condition can also be met in principle, since the specified gate sets are sufficient for universal MBQC [1, 43].
@@
To understand condition (2) in more detail, note that if the full graph state were built using HoS and two-qubit gates on the auxiliary spins, then each emitted photon would have a local Clifford error of the form $UZ^{m}$, where $U$ is a fixed Clifford and $m\in\{0,1\}$ is determined by the decoupling measurement on the emitter spin. If $M$ is the measurement to be subsequently performed on the photon in the MBQC protocol, then when correcting for the Clifford error, the effective measurement would be $M^{\prime}=UZ^{m}MZ^{m}U^{\dagger}$. When $M=Z$, then $M^{\prime}=UZU^{\dagger}$; or when $M=\cos\phi X+\sin\phi Y$, then $M^{\prime}=(-1)^{m}UZU^{\dagger}$. In the first case, the dependence on $m$ is completely removed, and the photon can be equivalently measured with $UZU^{\dagger}$ immediately after it is emitted in the HoD scheme. In the second case, it can also be immediately measured with $UZU^{\dagger}$, but now one must perform a bit flip on the classical measurement outcome if $m=1$; this is because an overall $-1$ factor on a spin observable simply flips the spin-up/spin-down outcomes. In summary, the correct computation can be attained through HoD and classical post-processing for observables of the specified form.
@@
While satisfying conditions (1) and (2) above is sufficient for universal MBQC using HoD, this is often not the most efficient method in terms of the overall number of photons used to drive the computation. For example, building out the graph to a certain depth and measuring in a different sequence than the emission order can lead to more compact gate implementations [43]. Also, using MBQC measurements outside of the x-y plane allows for more general forms of information flow [44]. Nevertheless, as we show in the next section, the HoD scheme can enable dramatically faster construction and higher fidelity with no photon storage required. For the HoS (or any non-destructive heralding) scheme, photon storage is required for at least the time to perform the spin-spin entangling gates and the decoupling measurement on the emitter.
Figure 2: Herald-on-detection (HoD) scheme. For certain MBQC protocols, it is sufficient to perform the deterministic entangling gate between the emitter and auxiliary spin only upon successful detection of the emitted photon. The photonic graph state constructed in this case is virtual (represented by dashed boundaries) and exists as a set of conditional phases stored on the auxiliary spin(s).
IV Scaling and fidelity for heralded schemes
A major benefit of both schemes comes in scaling up the size of graph states using near term hardware. The average time to successfully generate an $n_{p}$-photon graph by directly collecting photons from an emitter in $n_{p}$ subsequent excitation events for computation is $\mathcal{O}(\eta_{e}^{-n_{p}})$, where $\eta_{e}$ is the emitter collection efficiency, as any failed detection event truncates the graph. In typical deterministic quantum emitter-based schemes, any inefficiency in collection therefore leads to exponentially bad scaling with graph size. Given current hardware, this severely limits the size and rate of generation for photonic graph states, particularly those that require multiple spin qubits. We note that there are loss tolerance and percolation thresholds that improve this scaling, but they require much better collection efficiency than is feasible in near-term systems [45, 46, 47, 48]. For the heralded schemes, any failed detection of the emitter photon simply results in the reinitialization of the emitter and another attempt at the joint or single photon detection, while the graph under construction remains unaffected. Therefore, an $n_{p}$-photon graph state will be created over a time that is $\mathcal{O}(n_{p}\eta_{e}^{-1})$. As a comparison, the time to make an $n_{p}$-photon graph state (in units of the repetition period) is shown in Figure 3 for $\eta_{e}\in\{0.1,0.25,0.4\}$ for the deterministic (red), HoS (blue), and HoD (yellow) schemes. We do not include the time to perform gates and measurements on the spins for these estimates, and we assume that the emitter coherence does not limit the graph size for the deterministic scheme ($n_{p}/\tau R_{\rm{rep}}\ll 1$ where $\tau$ is the spin coherence time and $R_{\text{rep}}$ is the repetition rate of excitation). In practice, however, the short coherence times of the most efficient photon emitters further limit the size of graphs that can be produced in the deterministic scheme [32, 10]. For the heralded schemes, all auxiliary spins must remain coherent for the entire time they are a part of the graph, which is much longer than the $n_{p}$ repetition cycles over which the graph is generated in the deterministic scheme. Thus, moving from the deterministic scheme to one proposed here effectively means moving from a scheme limited by the photon collection efficiency to a scheme limited by spin coherence. If a hybrid of the HoS and HoD schemes is employed, the generation time falls somewhere between the curves for those schemes, with the fraction of photons generated via each method determining exactly where.
@@ IV Scaling and fidelity for heralded schemes
Figure 3: Time to make photonic graph states of size $n_{p}$ in units of the repetition period. The deterministic (red), HoS (blue), and HoD (yellow) schemes are compared. The three curves of each color represent emitter photon collection efficiencies $\eta_{e}\in\{0.1,0.25,0.4\}$. Polynomial scaling in the heralded schemes allows for the construction of larger graphs. Results are omitted where decoherence and false heralds bring the fidelity under $50\%$ for reasonable experimental parameters. In the HoS scheme, fidelity can be sacrificed to improve the rate for the construction of small graph states if desired. A full description of these estimates is given in Appendix D.
For the two proposed schemes there are additional factors that affect the fidelity that are not present in the deterministic scheme, including decoherence of the emitter and auxiliary spin(s) during the longer construction time and (for the HoS scheme) multiple photon pair production. Here we investigate the scaling of these sources of infidelity and compare them to the limits set by imperfect two-qubit spin-spin entangling gates, as well as the initial entanglement fidelities of the emitter photon and photon pair sources. We require one extra spin-spin entangling gate each time we add a photon to the graph, in addition to the gates required to make the desired final graph [37], which exponentially suppresses the fidelity of large photonic graph states if the gates are not perfect. Additionally, any infidelity of the photon pair source in the HoS scheme adds to the overall infidelity. Here, we demonstrate that if sufficient error correction or entanglement purification is possible [49], the additional sources of infidelity inherent to our scheme do not themselves prohibit exceedingly large graph states. For a full discussion of our experimental model and relevant sources of infidelity, see Appendix D.
Figure 4: Graph state fidelity versus photon number ($n_{p}$) for several different regimes of operations, and variable emitter-photon collection efficiencies ($\eta_{e}$) and dephasing timescales ($t_{\text{rep}}/\tau$). Gate and pair generation fidelities (grey) exponentially preclude the construction of large graph states in any scheme utilizing our “emit then add” method. With sufficient error correction or purification around these fidelities, regimes limited by false heralds and decoherence (blue) and decoherence alone (yellow), allow for the construction of significantly larger graph states.
In modeling decoherence under the usual dephasing map, any emitters or auxiliary spins will remain invariant with a probability $D(t,\tau)=\tfrac{1}{2}(1+e^{-t/\tau})$, for the duration $t$ until they are projectively measured. In either scheme, the emitter is reinitialized with each attempt to add a new photon to the larger graph, and hence the time it is required to remain coherent is reset. Furthermore, the classical conditioning offered by this “emit then add” method means the emitter is only required to remain coherent for the attempts where the relevant joint or single photon measurements are a success. Any auxiliary spins, on the other hand, are required to remain coherent for the duration of time they remain a part of the larger graph state in construction. The average time for the successful addition of a new photon to the larger graph goes as $t_{\text{rep}}/P_{s}$, where $t_{\text{rep}}=R_{\text{rep}}^{-1}$ is the repetition period of the excitation, and $P_{s}$ is the probability of a successful joint or single photon measurement for the HoS or HoD schemes, respectively. Note that in the HoD scheme we assume $P_{s}=\eta_{e}$. Hence, the total contribution to the final state fidelity from the emitter and any auxiliary spins is
$F_{D}^{(e)}(n_{p})$ $=\left(\frac{1}{2}\left(1+e^{-t_{\text{rep}}/\tau}\right)\right)^{n_{p}},$ (1a)
$\langle F_{D}^{(s)}(n_{p},P_{s})\rangle$ $=\frac{1}{2}\left(1+e^{-n_{p}t_{\text{rep}}/P_{s}\tau}\right),$ (1b)

### v3

@@ Abstract
Quantum emitter-based schemes for the generation of photonic graph states offer a promising, resource efficient methodology for realizing distributed quantum computation and communication protocols on near-term hardware. We present a heralded scheme for making photonic graph states that is compatible with the typically poor photon collection from state-of-the-art coherent quantum emitters. We demonstrate that the construction time for large graph states can be polynomial in the photon collection efficiency, as compared to the exponential scaling of current emitter-based schemes, which assume deterministic photon collection. The additional overhead here consists of an extra spin qubit plus one additional spin-spin entangling gate per photon added to the graph. While the proposed scheme requires both non-demolition measurement and efficient storage of photons in order to generate graph states for arbitrary applications, we show that many useful tasks, including measurement-based quantum computation, can be implemented without these requirements. As a use-case of our scheme, we construct a protocol for secure two-party computation that can be implemented efficiently on current hardware. Estimates of the fidelity to produce graph states used in the computation are given assuming current and near-term fidelities for highly coherent quantum emitters.
@@ I Introduction
The vast promise of quantum information technologies for faster and more secure computational and information systems relies on entanglement as a primary resource. Traditional gate-based quantum computing requires the ability to perform sequences of joint operations across multiple qubits. An alternative paradigm, known as measurement-based quantum computation (MBQC), realizes universal computation through sequences of single qubit measurements made on an initially prepared entangled resource state [1]. This is an appealing approach for photonic quantum systems as single photon rotations and measurements are straightforward using commercial optical elements, and fast efficient routing solutions are readily available [2]. Furthermore, the sequential nature of photon emission allows entangled states to be built from photons emitted at different times by the same emitter [3]. By using entangled emitters, it then becomes possible to simulate entangling gates between photons and overcome the difficulty of realizing direct photon-photon interactions [4]. Indeed, it is known that computationally useful graph states of photons can be generated using either a small number of coherent quantum emitters [5, 6], a combination of a single emitter and fusion gates [7, 8, 9], or a single quantum emitter in a feedback scheme [10].
@@
All of these aforementioned schemes assume a photon is successfully added to the graph every time an emitter is excited, and we thus term this class of schemes as being “deterministic”. For realistic systems with less than perfect emission and collection efficiency, the time to make a graph thus scales exponentially in the size of the graph because any failure to detect a photon triggers a restart of the whole protocol. In general, highly efficient collection from individual quantum emitters is a challenge that remains largely out of reach today, making such deterministic schemes impractical for generating even moderately sized graph states ($10-100$ photons) on near-term hardware [11, 12, 13, 14].
We introduce in Sec. II a method for photonic graph generation based on coherent emitters that uses an “emit-then-add” approach, where each photon is attached to the graph only following confirmation of its emission and collection. In this way, a lost photon does not truncate the rest of the graph and the time to produce a photonic graph state scales polynomially in the size of the graph. Compared to the deterministic protocols, one additional spin is required in order to allow the emitter to be disconnected from the graph until photon collection is confirmed. Furthermore, for each photon added to the graph, our scheme requires one additional spin-spin entangling operation plus one mid-circuit measurement and reset (MCMR) of the emitting spin. (Note that these only occur upon heralding of successful photon emission, they are not required on every attempt.)
Typically, confirming the presence of a photon in a particular optical mode requires destructive measurement of the photon. For arbitrary applications of photonic graph states where the photons are measured in any order and in any basis, one would thus require quantum non-demolition (QND) measurement [15, 16, 17], in order to determine successful collection of the photon before adding it to the larger graph, plus storage of the photon until it can be measured in the desired basis. Such a scheme is largely out of reach today, though we discuss in Appendix C a method for implementing this on current hardware based on entanglement swapping with a separately produced entangled photon pair.
Importantly, we show here that a large class of graph state protocols (including MBQC) admit projective measurements of each photon before adding it to the larger graph. In this way we show how to build a heralded virtual graph state of photons that never exist at the same time. Large virtual graph states can be built quickly with high fidelity in this way. We introduce an example use case in Sec. III, a particular protocol that uses multiple copies of a small virtual graph state to implement secure two-party computation on arbitrarily large inputs and requires just a single quantum emitter and two auxiliary spin qubits. We demonstrate how the coherence and entanglement fidelity requirements for this protocol are likely met by current state-of-the-art to near-term trapped ion and trapped neutral atom systems.
II Emit-then-add
Given the generally poor collection efficiency shown by highly coherent quantum emitters that can be entangled with additional spin qubits we propose a general methodology, dubbed “emit-then-add,” for constructing photonic graph states to circumvent this issue. A single emitter is initialized into an unentangled state and excited to produce a single photon that is entangled with its long-lived internal spin state and collected with some overall efficiency $\eta_{e}$. This can be implemented with a wide variety of quantum emitters including laser-cooled atoms or ions, quantum dots, and defects or dopants in wide-bandgap semiconductors [3, 18, 19, 14, 20, 21]. Photons can be encoded in various degrees of freedom including polarization, time-bin, and frequency [22, 23, 24, 25, 26]. In addition to the single emitter, we require a set of auxiliary spin qubits that can be controllably entangled in a pairwise way (with the emitter and with each other) via local and deterministic two-qubit spin-spin entangling gates. We note that these auxiliary spins are never used for emission of photons for the graph and can be different physical qubits from the emitter (i.e., a different atomic species in an atom or ion system or nearby nuclear spins coupled to a defect or dopant emitter in solid-state) [27, 28, 29]. We see that our protocol works best for an emitter optimized for collection efficiency and fast and high-fidelity preparation and readout, while the auxiliary spins are optimized for coherence.
@@ II Emit-then-add
Deterministic schemes for constructing photonic graph states, such as those in [3, 6], center their building operations on the emitting spin itself, while the central feature of our method is to add a photon to the graph only following a logical herald of its successful emission and collection. Successfully adding each photon to the graph state requires entangling the emitter with an auxiliary spin, and then measuring it out of the graph and reinitializing it to prepare for the next attempted photon emission. Following a correction to the photon and auxiliary spin based on the outcome of this measurement, the state of the system is as if the auxiliary spin had directly emitted the photon itself. In this sense, we construct a graph state of only heralded photons, where we have split the role of the emitter in deterministic schemes into an emitting spin that generates the photons and an auxiliary spin that stores the quantum information. Thus, we see that emit-then-add admits the construction of arbitrary photonic graph states using the same methods described in [6] with minimal additional overhead. Namely, the graph state itself is built using a sequence of two-qubit and local Clifford gates on the spins and only local Clifford gates on the photons. The local gates on each photon can all be combined and applied as a single rotation prior to using the graph state in subsequent applications.
The key improvement of our scheme is that uncollected photons do not disturb the state of the graph that is being built. When successive photon collection is required for building a graph on a single emitter, any failed detection traces the photon out of the system and truncates the graph, which necessitates restarting construction from scratch. In typical deterministic quantum emitter-based schemes, any inefficiency in collection therefore leads to exponentially poor scaling with graph size. Given current hardware, this severely limits the size and rate of generation for photonic graph states, particularly those that require multiple spin qubits. We note that there are loss tolerance and percolation thresholds that improve this scaling, but they require much better collection efficiency than is feasible with near-term systems [30, 31, 32, 33, 34, 35]. In our scheme, any failed detection simply results in the reinitialization of the emitting spin without any disturbance to the overall graph under construction.
For general purpose applications, we require a method to detect the presence of a photon that is non-destructive. This can be met with any form of QND measurement that heralds spin-photon entanglement. In Appendix C we provide an example for implementing such a scheme on current hardware, utilizing entanglement swapping and a separately produced entangled photon pair. Here, a pair of entangled photons is probabilistically produced, which can be implemented via standard nonlinear optical processes, such as spontaneous parametric down conversion (SPDC) [36, 37]. One member of the photon pair, the signal photon, is wavelength and bandwidth matched to the emitter photon, and the entangled pair is encoded in the same degree of freedom as the emitter photon. The emitter photon and signal photon are sent to a joint measurement apparatus which, upon a successful measurement outcome, projects the emitter spin and the unmeasured photon, the idler photon, onto an entangled Bell state [38]. Since only certain measurement outcomes correspond to entanglement between the emitter spin and idler photon, the procedure is repeated until a successful herald is flagged. While this procedure works in principle, it has some major drawbacks including a significant reduction in rate due to the probabilistic nature of generating entangled photon pairs as well as the introduction of additional infidelities due to factors including multi-pair production and imperfect heralding efficiency.
Figure 1: A construction scheme for heralding virtual graph states, with emit-then-add. A quantum emitter (red) is repeatedly excited until a single photon (orange) is successfully detected. Upon detection, logic is executed which passes the conditional phase information written on to the emitter to an auxiliary memory spin (black), which encodes a virtual graph state (represent by dashed boundaries).
@@
Instead, we show that for many applications, including MBQC, non-destructive measurement is not required. For these applications the nodes in the graph are measured sequentially with the choice of measurement basis on one node depending on outcomes of previous ones. It is, thus, not necessary to build the full graph state before beginning these measurements, as an emitter photon can be destructively measured as soon as the correct basis for measurement is determined [39]. This measurement thus serves doubly as a prescribed step in the MBQC protocol and as a herald for the emitter photon. Upon failure to detect the photon, we reinitialize the emitter and attempt generation again, as we would with some QND measurement. Upon successful detection, we perform all required logic on the emitter and auxiliary spins, measure the emitter out of the graph, and reinitialize the emitter to attempt generation of the next photon. This simpler emit-then-add scheme is shown in Fig. 1, and it has the advantage that it requires no storage of the emitter photon prior to measurement. The graph state being constructed is therefore “virtual” in the sense that not all photons within the graph need exist at the same time.
@@
There are limitations, however, for when this scheme can be employed. Namely, a photon can be destructively measured immediately upon generation provided that two conditions are satisfied: (1) the correct measurement basis for that photon, as set by the protocol, is determined prior to its emission, and (2) this measurement is either Pauli $Z$ or of the form $\cos\phi X+\sin\phi Y$. Condition (1) can be met for MBQC, as the emission order can be chosen to match the measurement order of the computation, albeit at the cost of using extra auxiliary spins and two-qubit gates in some cases [40]. The second condition can also be met in principle, since the specified gate sets are sufficient for universal MBQC [1, 41].
@@
To understand condition (2) in more detail, note that if the full graph state were built such that the emitter is measured before its emitted photon, then each photon would have a local Clifford error of the form $UZ^{m}$, where $U$ is a fixed Clifford and $m\in\{0,1\}$ is determined by the decoupling measurement on the emitter spin. If $M$ is the measurement to be subsequently performed on the photon in the MBQC protocol, then when correcting for the Clifford error, the effective measurement would be $M^{\prime}=UZ^{m}MZ^{m}U^{\dagger}$. When $M=Z$, then $M^{\prime}=UMU^{\dagger}$; or when $M=\cos\phi X+\sin\phi Y$, then $M^{\prime}=(-1)^{m}UMU^{\dagger}$. In the first case, the dependence on $m$ is completely removed, and the photon can be equivalently measured with $UMU^{\dagger}$ immediately after it is emitted in this simpler scheme. In the second case, it can also be immediately measured with $UMU^{\dagger}$, but now one must perform a bit flip on the classical measurement outcome if $m=1$; this is because an overall $-1$ factor on a spin observable simply flips the spin-up/spin-down outcomes. In total, the correct computation can still be attained, because protocols admitting only measurements of this form write their outcomes as a conditional phase on their neighbors in the graph. In our scheme, this conditional phase information is imparted on the emitter with each successful single photon detection, and subsequently passed into memory on one of the auxiliary spins. Computation is then achieved by passing conditional phases into memory in a specific pattern, performing logic between auxiliary spins, and classical post-processing for observables of the specified form.
@@
While satisfying conditions (1) and (2) above is sufficient for universal MBQC using virtual graph states, this is often not the most efficient method in terms of the overall number of photons used to drive the computation. For example, building out the graph to a certain depth and measuring in a different sequence than the emission order can lead to more compact gate implementations [41]. Also, using MBQC measurements outside of the x-y plane allows for more general forms of information flow [42]. Nevertheless, as we show below, this simpler scheme can enable dramatically faster construction and higher fidelity with no photon storage required. For emit-then-add schemes incorporating some form of QND measurement instead, photon storage is required for at least the time to perform the spin-spin entangling gates and the decoupling measurement on the emitter. For the remainder of this section, we discuss in more detail the advantages of emit-then-add, specifically in the context of this simplified construction scheme for heralding virtual graph states that contain $n_{p}$ photons.
@@
Figure 2: Time to make photonic graph states of size $n_{p}$ in units of the repetition period. Principal deterministic schemes (red) and our emit-then-add (blue) scheme are compared. The three curves of each color represent emitter photon collection efficiencies $\eta_{e}\in\{0.1,0.5,0.9\}$. The polynomial scaling in our schemes allows for the construction of larger photonic graphs. Cutoffs (grey, vertical) are depicted where the additional infidelity associated with emit-then-add precludes the construction of larger graph states with fidelities better than $1/2$, for $F_{\text{add}}\in\{0.99,0.999,0.9999\}$. Despite these cutoffs, graphs states of $10-100$ photons are achievable on realistic timescales, with only moderate requirements on $F_{\text{add}}$.
Overhead. It is known that the sequential nature of photon emission imposes nontrivial resource requirements when constructing graph states, specifically in terms of the number of required emitters and two-qubit spin-spin entangling gates needed to build out the graph [40]. As discussed in the introduction, constructing graph states with emit-then-add further adds to this overhead. The total number of spins we require is only one more than the deterministic case, as each auxiliary spin in our scheme replaces an emitter in deterministic schemes and only a single emitting spin is required. In practice, more emitting spins can be employed in parallel to speed up the construction. As the emitting spin in our scheme must remain disjoint from the graph for each excitation, an additional two-qubit spin-spin entangling gate is required per successfully heralded virtual photon, in order to entangle the emitter with the larger graph. In Appendix B, we demonstrate efficient construction subroutines for building certain graph states that directly employ the methods of [6].
Fidelity. In addition to overhead, there are infidelities that affect the final graph state with emit-then-add that are not present in the deterministic scheme, namely any infidelity in the additional two-qubit spin-spin entangling gate and mid-circuit measurement and reset of the emitter. We capture all of this, plus the initial fidelity of the emitter-photon entanglement, in a parameter, $F_{\text{add}}$, for the total fidelity of the process to add a photon to the graph. Note that photon detection inefficiency does not affect the fidelity because failing to detect a photon simply triggers a reset. Any infidelity in measuring the state of the photon does affect $F_{\text{add}}$, but this can be minimal for photons (and has the same effect as for the deterministic schemes for photonic graph state production).
Importantly for our scheme, when building a graph all auxiliary spins must remain coherent for the entire time they are a part of the graph, which is much longer than the $n_{p}$ repetition cycles over which the graph is generated in the deterministic scheme. Thus, moving from the deterministic scheme to one proposed here effectively means moving from a scheme limited by the photon collection efficiency to a scheme limited by spin coherence (denoted $\tau$). In modeling the effects of this dephasing, we note that the emitter is reinitialized with each attempt to add a new photon to the larger graph, and hence it is only required to remain coherent for time to add it to the larger graph, denoted $t_{\text{add}}$. For the auxiliary spins, on the other hand, the average time for the successful addition of a new photon to the larger graph goes as $t_{\text{rep}}/\eta_{e}$ where $t_{\text{rep}}^{-1}$ is the experimental repetition rate. Hence, the total contribution to the final state fidelity from the emitter and any auxiliary spins is
$F_{D}^{(e)}(n_{p})$ $=\left(\frac{1}{2}\left(1+e^{-(t_{\text{rep}}+t_{\text{add}})/\tau}\right)\right)^{n_{p}},$ (1a)
$\langle F_{D}^{(s)}(n_{p})\rangle$ $=\frac{1}{2}\left(1+e^{-n_{p}(t_{\text{rep}}/\eta_{e}+t_{\text{add}})/\tau}\right),$ (1b)